# A Systematic Search of Distant Superclusters with the Subaru Hyper Suprime-Cam Survey

### v1

@@ 4 CAMIRA supercluster catalog
Figure 9: Left panel: HSC $riz$ color composite image of the supercluster candidate with the highest multiplicity (5), at redshift $z\approx 0.87$, in our catalog. The field of view is $16^{\prime}\times 36^{\prime}$. White contours represent galaxy number density, stepping three galaxies/pixel from one galaxy/pixel. Right panel: zoom-in HSC $riz$ composite image of each member cluster. The field of view is $3^{\prime}\times 3^{\prime}$.
@@
Supercluster ID CAMIRA ID Multiplicity R.A. Dec. $z_{CAM}$ $z_{BCG}$ $N_{mem}$ Conf-z
(1) (2) (3) (4) (5) (6) (7) (8) (9)
1 J011124.2+014921 2 17.850688 1.822534 0.8077 0.81178 35.5 1
1 J011132.0+015238 2 17.883536 1.877400 0.8076 --1.00000 32.4 1
2 J013636.9+004213 3 24.153902 0.703794 0.7910 0.79443 38.4 1
2 J013708.9+002551 3 24.287286 0.431092 0.7918 0.79053 20.5 1
2 J013728.5+003623 3 24.368642 0.606420 0.7920 0.79216 36.0 1
3 J020832.7-043315 2 32.136452 --4.554213 0.7618 --1.00000 38.7 1
3 J020846.1-042624 2 32.192284 --4.440218 0.7580 --1.00000 42.0 1
4 J023527.9-010129 2 38.866387 --1.024991 0.7580 --1.00000 27.1 1
4 J023554.8-010029 2 38.978506 --1.008286 0.7532 --1.00000 27.5 1
5 J084911.7+024148 2 132.298617 2.696919 0.8292 --1.00000 33.6 1
5 J084941.2+025543 2 132.421754 2.928781 0.8304 --1.00000 26.6 1
6 J090755.5+025422 2 136.981335 2.906202 0.7989 0.79685 36.7 1
6 J090823.8+025132 2 137.099013 2.859133 0.8008 --1.00000 30.9 --1
7 J104319.9+041319 3 160.833029 4.222015 0.5700 0.57200 27.8 1
7 J104355.6+041313 3 160.981508 4.220447 0.5720 --1.00000 33.6 --1
7 J104401.7+040340 3 161.007286 4.061376 0.5700 0.57345 24.3 1
8 J104915.4+043307 2 162.314297 4.552191 0.7810 --1.00000 25.1 --1
8 J104926.6+043822 2 162.360939 4.639518 0.7802 --1.00000 39.0 1
9 J114545.0+015648 2 176.437664 1.946910 0.6060 0.63232 28.1 --1
9 J114559.6+020239 2 176.498319 2.044430 0.6040 0.63247 26.2 --1
10 J124250.2+035922 2 190.709164 3.989536 0.6458 --1.00000 28.1 1
10 J124312.6+035945 2 190.802338 3.996053 0.6503 --1.00000 25.9 --1
11 J134341.1+032930 3 205.921269 3.491779 0.7510 --1.00000 29.4 1
11 J134343.1+033630 3 205.929784 3.608402 0.7560 --1.00000 26.5 1
11 J134405.3+032517 3 206.021919 3.421467 0.7559 --1.00000 16.7 1
12 J142044.0+030004 4 215.183491 3.001134 0.8723 --1.00000 28.2 1
12 J142105.4+025611 4 215.272457 2.936596 0.8693 --1.00000 13.8 --1
12 J142114.8+030736 4 215.311775 3.126804 0.8789 --1.00000 26.6 1
12 J142141.9+030351 4 215.424449 3.064213 0.8820 --1.00000 12.2 --1
13 J145943.8+014807 2 224.932406 1.802151 0.5133 0.52117 25.4 1
13 J150032.1+015321 2 225.133663 1.889420 0.5250 0.52039 26.8 1
14 J152916.9+434057 4 232.320532 43.682659 0.6420 --1.00000 41.7 1
14 J152922.6+435005 4 232.344015 43.834781 0.6400 --1.00000 12.4 --1
14 J152942.6+433459 4 232.427692 43.583191 0.6280 0.64033 28.4 1
14 J152957.7+432446 4 232.490454 43.413040 0.6348 0.63663 45.1 1
15 J221359.5+015127 2 333.497897 1.857536 0.6860 0.68319 57.5 1
15 J221452.3+014439 2 333.717980 1.744227 0.6961 0.68270 46.0 1
16 J230235.1+000234 3 345.646100 0.042840 0.5155 0.51647 78.8 1
16 J230241.7-000830 3 345.673778 --0.141791 0.5141 0.51937 15.7 1
16 J230333.9+000834 3 345.891215 0.143036 0.5020 0.51599 54.9 1
17 J230545.0+031154 3 346.437649 3.198562 0.9340 --1.00000 10.1 --1
17 J230550.4+031831 3 346.460123 3.308698 0.9300 --1.00000 26.4 --1
17 J230632.6+030754 3 346.635948 3.131718 0.9285 --1.00000 26.3 --1
18 J232840.3+043506 2 352.168095 4.585250 0.6300 --1.00000 27.4 --1
18 J232905.3+042955 2 352.271918 4.498687 0.6260 --1.00000 28.2 --1
@@
Table 1: The first 18 supercluster candidates in the CAMIRA supercluster catalog. The columns are: (1) Supercluster ID (2) Cluster ID (Cluster ID in the released catalog will have a prefix “HSCCL”, indicating the object type) (3) Multiplicity of superclusters (4) Right ascension [degree, J2000] (5) Declination [degree, J2000] (6) Photometric redshift of the CAMIRA cluster (7) Spectroscopic redshift of the BCG if available (8) Richness (9) The confident redshift flag. It is $1$ if the photometric redshift of the cluster meets one of the criteria described in Section 2, otherwise it is $-1$. The complete catalog is available online.
@@
The final catalog consists of 633 supercluster candidates between redshift 0.5 and 1.0. It is presented in Table 1. Figure 9 shows the HSC $riz$ composite image of the supercluster candidate with the highest multiplicity in the final catalog. We believe that this is the first supercluster catalog that extends beyond redshift $z=0.7$. However, since more than half of the clusters do not have confident redshifts, we must carefully examine the impact of photo-$z$ uncertainty and conservatively call them “supercluster candidates” (Section 4.1). We also examine the multiplicity function (Section 4.2) and total mass of our supercluster candidates (Section 4.3). Finally, we compare the supercluster candidates that we detect with two known high-$z$ superclusters (Section 4.4).
@@ 4.1 Effect of photometric redshift uncertainty
Taking into account the photo-$z$ uncertainty, we next estimate the contamination rate of the supercluster catalog. To do so, we compare the supercluster catalogs obtained from the unperturbed light cones with those from the perturbed light cones. Assuming the supercluster catalogs extracted from the unperturbed light cones as standard, we define the contamination rate as the ratio between the number of perturbed superclusters not belonging to any unperturbed superclusters and the number of perturbed superclusters. Figure 12 shows the contamination rate as a function of redshift. The contamination rate increases from around $0.15$ at $z\approx 0.5$ to $0.50$ at $z\approx 1.0$ due to the increasing perturbed fraction in Figure 3.
@@
We regard a supercluster candidate as confirmed if all of its member clusters possess a confident redshift (see Figure 3). 164 superclusters are confirmed among the 633 supercluster candidates, constituting one-fourth of the catalog presented here. However, we note that this does not imply that the remaining supercluster candidates are spurious. This is because the three criteria described in Section 2 can not falsify the photo-$z$ of a cluster. A cluster photo-$z$ may not satisfy the criteria because of the incompleteness of the spectroscopic survey, the limited coverage of DESI EDR or the greater photo-$z$ scatter of DEmP.
@@ 4.2 Multiplicity Function
Refer to captionRefer to caption
Figure 13: The multiplicity function at redshift $z=[0.55,0.65]$ (left panel) and $z=[0.85,0.95]$ (right panel). The gray histograms are multiplicity functions of ground-truth superclusters. The green histograms and blue histograms are multiplicity functions of superclusters extracted from unperturbed light cones and perturbed light cones, respectively. The error bars here indicate the $1\sigma$ standard deviation in six realizations. The red histogram shows the multiplicity function of CAMIRA supercluster candidates. The consistency between multiplicity function of CAMIRA supercluster candidates and mock perturbed superclusters again suggests the impact of photo-$z$ uncertainty.
@@
The multiplicity is defined as the number of member clusters in a supercluster, and the multiplicity function describes the multiplicity distribution. We show the multiplicity function of ground-truth superclusters, supercluster candidates from unperturbed/perturbed light cones, and CAMIRA supercluster candidates in Figure 13. The multiplicity function of ground-truth superclusters declines rapidly as the multiplicity increases. Superclusters having four members are ten times rarer than cluster pairs, and superclusters having six members are 100 times more scarce. This demonstrates that the supercluster population is dominated by cluster pairs and triplets, suggesting the limit of gravitationally bounded structures. The multiplicity functions extracted from unperturbed light cones well recover the ground-truth multiplicity functions by our method, at least below multiplicity of eight. We attribute the absence of high-multiplicity superclusters in unperturbed light cones to the following reasons: (1) Our optimization strategy prioritizes purity over completeness as described in Section 3.3, which prefers a shorter linking length due to the purity-completeness trade-off. (2) Although we weigh each supercluster by its mass in the performance metrics, the contribution of these high-multiplicity superclusters is still low, since they are extremely rare. (3) The existence of these high-multiplicity superclusters is extremely rare and stochastic, as revealed by the large error bars. We note that we do recover high-multiplicity superclusters in intermediate redshift bins not shown here. Capturing these scarce high-multiplicity superclusters will be as a major goal for improvement of our method in the future.
@@ 4.4.1 King Ghidorah Supercluster
Figure 15: CAMIRA supercluster candidates and isolated clusters in King Ghidorah Supercluster field overlaid by HSC PDR3 density map in two different redshift slices. The orange open circles represent the positions of CAMIRA clusters that are not associated with any supercluster candidates. Red crosses and red squares indicate the positions of supercluster candidates with multiplicity two ($m=2$) and four ($m=4$), respectively (here the positions are the mean positions of member clusters). The number labeled under each supercluster candidate is the mean redshift of member clusters. If all members of a given supercluster have confident redshift (see Section 2.3), then the redshift number is colored in green; otherwise redshift number is colored in yellow. Under the definition that superclusters will eventually experience gravitational collapse, our results suggest that KGSC might be fragmented into numerous supercluster candidates.
@@
Figure 15 depicts the positions of the supercluster candidates of different multiplicities along with the CAMIRA clusters that are not associated with any supercluster candidates in the KGSC field. Under the definition that superclusters will eventually experience gravitational collapse, KGSC is fragmented into numerous supercluster candidates. Three supercluster candidates are associated with the density peak in the northern main body, while the rest of the CAMIRA clusters may eventually evolve into island universes. Although we find two supercluster candidates connected with the northeast filaments, we have less confidence in the photometric redshift of these cluster members. Interestingly, there appear to be at least three supercluster candidates associated with the southeast filaments, including one supercluster candidate with four confident redshift cluster members. Moreover, there is another supercluster candidate with two members residing close to the four-member supercluster candidate, indicating an extremely dense environment here.
@@ 4.4.2 CL1604
Figure 16 shows the CAMIRA clusters that are associated with supercluster candidates and the isolated clusters with HSC PDR3 density map (Shimakawa et al., 2021) overlaid. Similarly to the case of the KGSC, we find that only parts of the structures are identified as superclusters under the constraint of future collapse. In this case, there is one cluster pair at the core of the main body and another in the northeast extending structure, suggesting the rest of clusters associated with Cl1604 will evolve into island universes.
@@ 5 Clusters and Brightest Cluster Galaxies in superclusters
As a demonstration of the use of this supercluster catalog, in this section, we investigate the environmental effects of superclusters on their member clusters and BCGs. We refer to clusters that reside inside any supercluster candidates as supercluster clusters (SC-Cls), while we define our control sample, isolated clusters (Iso-Cls), as clusters that are not associated with any supercluster candidates. We further require that the clusters in Iso-Cls meet the following two criteria to avoid any potential associations with supercluster candidates: (1) no other clusters are located within a projected distance of $10\,h^{-1}\,$cMpc (2) no other clusters are within the line-of-sight distance ±$\pm 0.03(1+z_{cl})$. This set of samples is referred to as All.
@@
Taking into account the possible impact of the photo-$z$ uncertainty as demonstrated in Section 4.1, we construct another set of SC-Cls and Iso-Cls using only clusters with confident redshift described in Section 2.3 and Figure 3. For this second sample, we regard a supercluster candidate to be confirmed if all of its member clusters have a confident redshift. For SC-Cl, we consider clusters from confirmed superclusters. For the control sample, we include clusters with confident redshift that are not associated with any confirmed superclusters. We also require that clusters in Iso-Cls meet the two criteria mentioned above. This second sample is referred to as Conf-z. Moreover, we further split each set of samples into two redshift bins. Table 3 summarizes the number of our SC-Cls and Iso-Cls in each redshift bin. For our BCG analysis, we determine the BCG for each CAMIRA cluster following Dalal et al. (2021), who selected the most massive (red) galaxy in a cluster with membership probability $w>0.1$.

### v2

@@ 4 CAMIRA supercluster catalog
Figure 9: Left panel: HSC $riz$ color composite image of one of the supercluster candidates with the highest multiplicity (5), at redshift $z\approx 0.87$, in our catalog. The field of view is $16^{\prime}\times 36^{\prime}$. White contours represent galaxy number density, stepping three galaxies/pixel from one galaxy/pixel. Right panel: zoom-in HSC $riz$ composite image of each member cluster. The field of view is $3^{\prime}\times 3^{\prime}$.
@@
Supercluster ID CAMIRA ID Multiplicity R.A. Dec. $z_{CAM}$ $z_{BCG}$ $z_{cl,DESI}$ $N_{mem}$ Conf-z
(1) (2) (3) (4) (5) (6) (7) (8) (9) (10)
1 J011124.2+014921 2 17.850688 1.822534 0.8077 0.81178 --1.00000 35.5 1
1 J011132.0+015238 2 17.883536 1.877400 0.8076 --1.$0$-1.00000 32.4 1
2 J013636.9+004213 3 24.153902 0.703794 0.7910 0.79443 --1.00000 38.4 1
2 J013708.9+002551 3 24.287286 0.431092 0.7918 0.79053 --1.00000 20.5 1
2 J013728.5+003623 3 24.368642 0.606420 0.7920 0.79216 --1.00000 36.0 1
3 J020832.7-043315 2 32.136452 --4.554213 0.7618 --1.$0$-1.00000 38.7 1
3 J020846.1-042624 2 32.192284 --4.440218 0.7580 --1.$0$-1.00000 42.0 1
4 J023527.9-010129 2 38.866387 --1.024991 0.7580 --1.$0$-1.00000 27.1 1
4 J023554.8-010029 2 38.978506 --1.008286 0.7532 --1.$0$-1.00000 27.5 1
5 J084911.7+024148 2 132.298617 2.696919 0.8292 --1.$0$-1.00000 33.6 1
5 J084941.2+025543 2 132.421754 2.928781 0.8304 --1.$0$-1.00000 26.6 1
6 J090755.5+025422 2 136.981335 2.906202 0.7989 0.79685 --1.00000 36.7 1
6 J090823.8+025132 2 137.099013 2.859133 0.8008 --1.$0$-1.00000 30.9 1
7 J095611.7+015934 2 149.048780 1.992959 0.9954 --1.00000 0.96491 49.1 1
7 J095624.3+015558 2 149.101237 1.932960 0.9700 --1.00000 0.96481 26.5 1
8 J104319.9+041319 3 160.833029 4.222015 0.5700 0.57200 --1.00000 27.8 1
8 J104355.6+041313 3 160.981508 4.220447 0.5720 --1.$0$-1.00000 33.6 1
8 J104401.7+040340 3 161.007286 4.061376 0.5700 0.57345 --1.00000 24.3 1
9 J104915.4+043307 2 162.314297 4.552191 0.7810 --1.$0$-1.00000 25.1 1
9 J104926.6+043822 2 162.360939 4.639518 0.7802 --1.$0$-1.00000 39.0 1
10 J114545.0+015648 2 176.437664 1.946910 0.6060 0.63232 --1.00000 28.1 1
10 J114559.6+020239 2 176.498319 2.044430 0.6040 0.63247 --1.00000 26.2 1
11 J115658.8+012525 3 179.245111 1.423803 0.5820 0.57220 0.56916 44.6 1
11 J115724.7+013829 3 179.352939 1.641495 0.5736 --1.00000 0.56814 12.7 1
11 J115745.9+012736 3 179.441162 1.460048 0.5539 0.56218 0.56827 27.7 1
12 J124250.2+035922 2 190.709164 3.989536 0.6458 --1.$0$-1.00000 28.1 1
12 J124312.6+035945 2 190.802338 3.996053 0.6503 --1.$0$-1.00000 25.9 1
13 J134341.1+032930 3 205.921269 3.491779 0.7510 --1.$0$-1.00000 29.4 1
13 J134343.1+033630 3 205.929784 3.608402 0.7560 --1.$0$-1.00000 26.5 1
13 J134405.3+032517 3 206.021919 3.421467 0.7559 --1.$0$-1.00000 16.7 1
14 J141026.0-011432 2 212.608325 --1.242480 0.6122 0.63922 0.64088 31.1 1
14 J141050.4-010930 2 212.710020 --1.158503 0.6300 0.63465 0.63915 34.6 1
15 J142044.0+030004 4 215.183491 3.001134 0.8723 --1.$0$-1.00000 28.2 1
15 J142105.4+025611 4 215.272457 2.936596 0.8693 --1.$0$-1.00000 13.8 1
15 J142114.8+030736 4 215.311775 3.126804 0.8789 --1.$0$-1.00000 26.6 1
15 J142141.9+030351 4 215.424449 3.064213 0.8820 --1.$0$-1.00000 12.2 1
16 J145943.8+014807 2 224.932406 1.802151 0.5133 0.52117 --1.00000 25.4 1
16 J150032.1+015321 2 225.133663 1.889420 0.5250 0.52039 --1.00000 26.8 1
17 J152916.9+434057 4 232.320532 43.682659 0.6420 --1.$0$-1.00000 41.7 1
17 J152922.6+435005 4 232.344015 43.834781 0.6400 --1.$0$-1.00000 12.4 --1
17 J152942.6+433459 4 232.427692 43.583191 0.6280 0.64033 --1.00000 28.4 1
17 J152957.7+432446 4 232.490454 43.413040 0.6348 0.63663 --1.00000 45.1 1
18 J221359.5+015127 2 333.497897 1.857536 0.6860 0.68319 --1.00000 57.5 1
18 J221452.3+014439 2 333.717980 1.744227 0.6961 0.68270 --1.00000 46.0 1
@@
Table 1: The first 18 supercluster candidates in the CAMIRA supercluster catalog. The columns are: (1) Supercluster ID (2) Cluster ID (Cluster ID in the released catalog will have a prefix “HSCCL”, indicating the object type) (3) Multiplicity of superclusters (4) Right ascension [degree, J2000] (5) Declination [degree, J2000] (6) Photometric redshift of the CAMIRA cluster (7) Spectroscopic redshift of the BCG if available (8) Cluster redshift derived using DESI EDR data, see Section 2.2 for detail (9) Richness (10) The confident redshift flag. It is $1$ if the photometric redshift of the cluster meets one of the criteria described in Section 2, otherwise it is $-1$. The complete catalog is available online.
@@
The final catalog consists of 673 supercluster candidates between redshift 0.5 and 1.0. It is presented in Table 1. Figure 9 shows the HSC $riz$ composite image of one of the supercluster candidates with the highest multiplicity in the final catalog. We believe that this is the first supercluster catalog that extends beyond redshift $z=0.7$. However, since approximately half of the clusters do not have confident redshifts, we must carefully examine the impact of photo-$z$ uncertainty and conservatively call them “supercluster candidates” (Section 4.1). We also examine the multiplicity function (Section 4.2) and total mass of our supercluster candidates (Section 4.3). Finally, we compare the supercluster candidates that we detect with two known high-$z$ superclusters (Section 4.4).
@@ 4.1 Effect of photometric redshift uncertainty
Taking into account the photo-$z$ uncertainty, we next estimate the contamination rate of the supercluster catalog. To do so, we compare the supercluster catalogs obtained from the unperturbed light cones with those from the perturbed light cones. Assuming the supercluster catalogs extracted from the unperturbed light cones as standard, we define the contamination rate as the ratio between the number of perturbed superclusters not belonging to any unperturbed superclusters and the number of perturbed superclusters. Figure 12 shows the contamination rate as a function of redshift. The contamination rate increases from around $0.05$ at $z\approx 0.5$ to $0.25$ at $z\approx 1.0$ due to the increasing perturbed fraction in Figure 3.
@@
We regard a supercluster candidate as confirmed if all of its member clusters possess a confident redshift (see Figure 3). 436 superclusters are confirmed among the 673 supercluster candidates, constituting $65\%$ of the catalog presented here. However, we note that this does not imply that the remaining supercluster candidates are spurious. This is because the three criteria described in Section 2 can not falsify the photo-$z$ of a cluster. A cluster photo-$z$ may not satisfy the criteria because of the incompleteness of the spectroscopic survey, the limited coverage of DESI EDR or the greater photo-$z$ scatter of DEmP.
@@ 4.2 Multiplicity Function
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 13: The multiplicity function at redshift $z=[0.55,0.65]$ (top left panel), $z=[0.65,0.75]$ (top right panel), $z=[0.75,0.85]$ (bottom left panel), and $z=[0.85,0.95]$ (bottom right panel). The gray histograms are multiplicity functions of ground-truth superclusters. The green histograms and blue histograms are multiplicity functions of superclusters extracted from unperturbed light cones and perturbed light cones, respectively. The error bars here indicate the $1\sigma$ standard deviation in six realizations. The red histogram shows the multiplicity function of CAMIRA supercluster candidates. The consistency between multiplicity function of CAMIRA supercluster candidates and mock perturbed superclusters again suggests the impact of photo-$z$ uncertainty.
@@
The multiplicity is defined as the number of member clusters in a supercluster, and the multiplicity function describes the multiplicity distribution. We show the multiplicity function of ground-truth superclusters, supercluster candidates from unperturbed/perturbed light cones, and CAMIRA supercluster candidates in Figure 13. The multiplicity function of ground-truth superclusters declines rapidly as the multiplicity increases. Superclusters having four members are ten times rarer than cluster pairs, and superclusters having six members are 100 times more scarce. This demonstrates that the supercluster population is dominated by cluster pairs and triplets, suggesting the limit of gravitationally bounded structures. The multiplicity functions extracted from unperturbed light cones well recover the ground-truth multiplicity functions by our method, at least below multiplicity of eight. We attribute the absence of high-multiplicity superclusters in unperturbed light cones to the following reasons: (1) Our optimization strategy prioritizes purity over completeness as described in Section 3.3, which prefers a shorter linking length due to the purity-completeness trade-off. (2) Although we weigh each supercluster by its mass in the performance metrics, the contribution of these high-multiplicity superclusters is still low, since they are extremely rare. (3) The existence of these high-multiplicity superclusters is extremely rare and stochastic, as revealed by the large error bars. (4) Some high-multiplicity superclusters are split or identified as low-multiplicity superclusters as we show in Appendix D. Capturing these scarce high-multiplicity superclusters will be as a major goal for improvement of our method in the future.
@@ 4.4.1 King Ghidorah Supercluster
Figure 15: CAMIRA supercluster candidates and isolated clusters in King Ghidorah Supercluster field overlaid by HSC PDR3 density map in two different redshift slices. The orange open circles represent the positions of CAMIRA clusters that are not associated with any supercluster candidates. Red crosses and red triangles indicate the positions of supercluster candidates with multiplicity two ($m=2$) and three ($m=3$), respectively (here the positions are the mean positions of member clusters). The number labeled under each supercluster candidate is the mean redshift of member clusters. If all members of a given supercluster have confident redshift (see Section 2.4), then the redshift number is colored in green; otherwise redshift number is colored in yellow. Under the definition that superclusters will eventually experience gravitational collapse, our results suggest that KGSC might be fragmented into numerous supercluster candidates.
@@
Figure 15 depicts the positions of the supercluster candidates of different multiplicities along with the CAMIRA clusters that are not associated with any supercluster candidates in the KGSC field. Under the definition that superclusters will eventually experience gravitational collapse, KGSC is fragmented into numerous supercluster candidates. Three supercluster candidates are associated with the density peak in the northern main body, while the rest of the CAMIRA clusters may eventually evolve into island universes. Although we find four supercluster candidates connected with the northeast filaments, two of them could not be confirmed. Interestingly, there appear to be at least three supercluster candidates associated with the southeast filaments, including one supercluster candidate with three confident redshift cluster members. This indicates an extremely dense environment here.
@@ 4.4.2 CL1604
Figure 16 shows the CAMIRA clusters that are associated with supercluster candidates and the isolated clusters with HSC PDR3 density map (Shimakawa et al., 2021) overlaid. Similarly to the case of the KGSC, we find that only parts of the structures are identified as superclusters under the constraint of future collapse. In this case, there is one cluster pair at the west of the main body and another in the northeast extending structure, suggesting the rest of clusters associated with Cl1604 will evolve into island universes.
@@ 5 Clusters and Brightest Cluster Galaxies in superclusters
As a demonstration of the use of this supercluster catalog, in this section, we investigate the environmental effects of superclusters on their member clusters and BCGs. Our cluster sample is divided into two redshift bins: $z=[0.5,0.7]$ and $z=[0.7,1.0]$. We refer to clusters that reside inside any supercluster candidates as supercluster clusters (SC-Cls), while we define our control sample, isolated clusters (Iso-Cls), as clusters that are not associated with any supercluster candidates. We further require that the clusters in Iso-Cls meet the following two criteria to avoid any potential associations with supercluster candidates: (1) no other clusters are located within a projected distance of $15\,h^{-1}\,$cMpc for lower redshift bin and $20\,h^{-1}\,$cMpc for higher redshift bin, (2) no other clusters are within the line-of-sight distance ±$\pm 0.03(1+z_{cl})$ for both redshift bins. This set of samples is referred to as All.
@@
Taking into account the possible impact of the photo-$z$ uncertainty as demonstrated in Section 4.1, we construct another set of SC-Cls and Iso-Cls using only clusters with confident redshift described in Section 2.4 and Figure 3. For this second sample, we regard a supercluster candidate to be confirmed if all of its member clusters have a confident redshift. For SC-Cl, we consider clusters from confirmed superclusters. For the control sample, we include clusters with confident redshift that are not associated with any confirmed superclusters. We also require that clusters in Iso-Cls meet the two criteria mentioned above. This second sample is referred to as Conf-z. Table 3 summarizes the number of our SC-Cls and Iso-Cls in each redshift bin. For our BCG analysis, we determine the BCG for each CAMIRA cluster following Dalal et al. (2021), who selected the most massive (red) galaxy in a cluster with membership probability $w>0.1$.